# From Hypothesis to Publication: A Comprehensive Survey of AI-Driven Research Support Systems

### v3

@@ Appendix D Ability Comparison
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2.5em, minimum width=5em, edge+=darkgray, line width=1pt, s sep=6pt, inner xsep=0.6em, inner ysep=0.9em, line width=0.8pt, text width=10em, ver/.style= rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=13em , leaf/.style= text opacity=1, fill opacity=.5, fill=hidden-blue!90, text=black, text width=50em, font=, inner xsep=0.6em, inner ysep=0.9em, draw, , , [ A survey of AI for Research, ver [ Hypothesis
Formulation  (§2) [ Knowledge Synthesize  (§2.1) [ Research Paper
Recommendation  (§2.1.1) [ PaperWeaver Lee et al. (2024), ArZiGo Pinedo et al. (2024), ComLittee Kang et al. (2023),  Stergiopoulos et al. (2024),  Kang et al. (2022),  Kreutz and Schenkel (2022) ,  Shahid et al. (2020),  Bai et al. (2019),  Li and Zou (2019),  Beel et al. (2016) , leaf, text width=50em ] ] [ Systematic
Literature Review  (§2.1.2) [ AutoSurvey  Wang et al. (2024e) , OpenScholar  Asai et al. (2024) , HiReview  Hu et al. (2024b), PaperQA2  Skarlinski et al. (2024), Knowledge Navigator  Katz et al. (2024), CHIME  Hsu et al. (2024), STORM  Shao et al. (2024), LitLLM  Agarwal et al. (2024a), Bio-SIEVE  Robinson et al. (2023), KGSum  Wang et al. (2022a), Agarwal et al. (2024b), Lai et al. (2024), Susnjak et al. (2024), Bolaños et al. (2024), Block and Kuckertz (2024), Zhu et al. (2023) , Altmami and Menai (2022) , leaf, text width=50em ] ] ] [ Hypotheses Generation  (§2.2) [ Other Works [ MOOSE-Chem  Yang et al. (2024c), IdeaSynth  Pu et al. (2024), Si et al. (2024), Kumar et al. (2024), Girotra et al. (2023), Qi et al. (2023), Krenn et al. (2022), Henry and McInnes (2017) , leaf, text width=50em ] ] [ Input Data Quality [ SciPIP Wang et al. (2024d), COI Li et al. (2024a), SciAgents Ghafarollahi and Buehler (2024), DATAVOYAGER  Majumder et al. (2024a), SciMON Wang et al. (2024c), Buehler (2024), Liu et al. (2024a) , leaf, text width=50em ] ] [ Hypothesis Quality [ Agentrxiv Schmidgall and Moor (2025) Dolphin Yuan et al. (2025), VIRSCI Su et al. (2024), Nova Hu et al. (2024a), AI Scientist Lu et al. (2024), SGA Ma et al. (2024), HypoGeniC  Zhou et al. (2024), ResearchAgent Baek et al. (2024), Acceleron Nigam et al. (2024), MOOSE  Yang et al. (2024b) , leaf, text width=50em ] ] ] ] [ Hypothesis Validation  (§3) [ Scientific Claim Verification  (§3.1) [ Claim [ HiSS Zhang and Gao (2023), SFAVEL Bazaga et al. (2024), ProToCo Zeng and Gao (2023), Wührl et al. (2024a) , leaf, text width=50em ] ] [ Evidence [ Glockner et al. (2024a), Vladika and Matthes (2024b), Vladika and Matthes (2024a), Wührl et al. (2024b), Pan et al. (2023b), Glockner et al. (2022), Wadden et al. (2022b) , leaf, text width=50em ] ] [ Verification [ ClaimVer Dammu et al. (2024), MAGIC Kao and Yen (2024), FactKG Kim et al. (2023), aedFaCT Altuncu et al. (2023), PROGRAMFC Pan et al. (2023a), Zhang et al. (2024b), Muharram and Purwarianti (2024), Eldifrawi et al. (2024), Cao et al. (2024b), Das et al. (2023), Wu et al. (2023) , leaf, text width=50em ] ] ] [ Theorem Proving  (§3.2) [ DeepSeek-Prover Xin et al. (2024), Lean-STaR Lin et al. (2024), Thor Jiang et al. (2022a), COPRA Thakur et al. (2024), Logo-power Wang et al. (2024b), Baldur First et al. (2023), Mustard Huang et al. (2024c), DT-Solver Wang et al. (2023b), HTPS Lample et al. (2022), GPT-f Polu and Sutskever (2020), Wang et al. (2024a), Song et al. (2024), Li et al. (2024e), Zhao et al. (2023b), Yang et al. (2023c), Jiang et al. (2023) , leaf, text width=62.5em ] ] [ Experiment Validation  (§3.3) [ MatPilot Ni et al. (2024), CRISPR-GPT Huang et al. (2024a), AgentHPO  Liu et al. (2024c), AutoML-GPT Zhang et al. (2023), MLCopilot Zhang et al. (2024a), Arlt et al. (2024), Kambhampati et al. (2024), Ruan et al. (2024b), Huang et al. (2024b), Boiko et al. (2023), Bran et al. (2024), Szymanski et al. (2023), Liu et al. (2023b), Manning et al. (2024), Mou et al. (2024) , leaf, text width=62.5em ] ] ] [ Manuscript
Publication  (§4) [ Manuscript Writing  (§4.1) [ Citation Text
Generation [ SciLit Gu and Hahnloser (2023), DisenCite Wang et al. (2022b), Gu and Hahnloser (2024), Gao et al. (2023b), Yu et al. (2022), Jung et al. (2022) , leaf, text width=50em ] ] [ Related Work
Generation [ ScholaCite Martin-Boyle et al. (2024), UR3WG Shi et al. (2023), Li and Ouyang (2024), Yu et al. (2024b), Nishimura et al. (2024), Li and Ouyang (2022) , leaf, text width=50em ] ] [ Complete Manuscripts
Generation [ Cocoa Feng et al. (2024), Step-Back Tang et al. (2024a), data-to-paper Ifargan et al. (2024), ARIES D’Arcy et al. (2024b), OREO Li et al. (2022), R3 Du et al. (2022a), Lai et al. (2024), Chamoun et al. (2024), Pividori and Greene (2024), Jourdan et al. (2023), Lin (2024) , leaf, text width=50em ] ] ] [ Peer Review  (§4.2) [ Paper Review
Generation [ CycleResearcher Weng et al. (2024), OpenReviewer Idahl and Ahmadi (2024) RelevAI-Reviewer Couto et al. (2024), AgentReview Jin et al. (2024), SWIF2T Chamoun et al. (2024), MARG D’Arcy et al. (2024a), Quality Assist Setio and Tsuchiya (2022), KID-Review Yuan and Liu (2022) Tan et al. (2024), Liang et al. (2023), Robertson (2023), Muangkammuen et al. (2022), , leaf, text width=50em ] ] [ Meta-Review
Generation [ Deeprview Zhu et al. (2025) PeerArg Sukpanichnant et al. (2024) GLIMPSE Darrin et al. (2024), MetaWriter Sun et al. (2024), CGI2 Zeng et al. (2023), Li et al. (2024b), Santu et al. (2024), Kumar et al. (2023), Li et al. (2023) , leaf, text width=50em ] ] ] ] ]

### v4

@@ Acknowledgments
Appendix A Further Discussion
We posit that AI for research constitutes a subset of AI for science. While AI for research primarily focuses on supporting or automating the research process, it is not domain-specific and places greater emphasis on methodological advancements. In contrast, AI for science extends beyond the research process to include result-oriented discovery processes within specific domains, such as materials design, drug discovery, biology, and the solution of partial differential equations (2310YizhenZhengLarge; 2311AI4ScienceThe; 2406YuZhangA).
Hypothesis generation, which is primarily based on literature-based review (LBD) (86DonRSwansonUndiscovered; 17SebastianEmerging), emphasizing the process by which researchers generate new concepts, solutions, or approaches through existing research and their own reasoning. Scientific discovery encompasses not only hypothesis generation, but also innovation in fields like molecular optimization and drug development  (2401GeyanYeDrugAssist; 24ShengchaoLiuConversational), driven by outcome-oriented results.
Existing research frequently addresses the systematic literature survey, which constitutes a component of the knowledge synthesis process during hypothesis formulation, alongside the related work generation phase in manuscript writing (25LuoZimingLLM4SR). However, we argue that these two tasks are distinct in nature. The systematic literature survey primarily focuses on summarizing knowledge extracted from diverse scientific documents, thereby assisting researchers in acquiring an initial understanding of a specific field (22NoufIbrahimAltmamiAutomatic). In contrast, related work generation focuses on the writing process, emphasizing selection of pertinent literature and effective content structuring  (24NishimuraToward).
  • In research paper recommendation, PaperWeaver (2403YoonjooLeePaperWeaver) offers an interactive page that allows users to modify the topics they are interested in.
  • In systematic literature review,  24BlockWhat highlights the significant role of humans, including setting correct questions and individualized problem-solving and theorizing. Meanwhile,  2407ChaoChunHsuCHIME emphasizes manual correction during the outline generation process.
  • In hypothesis generation, AI engages more closely with human researchers, ranging from scenarios where humans provide the core ideas and AI contributes by iteratively refining them  (2410KevinPuIdeaSynth), to more collaborative settings where humans and AI engage in dialogue to facilitate new scientific discoveries  (2411ZiqiNiMatPilot; 24ShengchaoLiuConversational; 2401GeyanYeDrugAssist).
  • In scientific claim verification,  2305EnesAltuncuaedFaCT; 2301AnubrataDasThe highlight the critical role of experts in countering fake scientific news and advocate for the incorporation of expert opinions as a form of evidence.
  • In theorem proving,  2404PeiyangSongTowards proposes leveraging LLMs as assistants to human researchers by generating suggested proof steps throughout the proving process.
  • In experiment Validation,  2411ZiqiNiMatPilot enhances the experimental setup through human-AI dialogue, whereas  2408RuochenLiMLRCopilot incorporates human input and real-time adjustments during the execution phase to optimize experimental design.
  • In manuscript writing,  2404TalIfarganAutonomous; 2412KJKevinFengCocoa; 2204WanyuDuRead require human intervention to suggest improvements to AI-generated paragraphs and enhance their quality through interactive methods.
  • In peer review,  2310SandeepKumarWhen; 2406MaximeDarrinGLIMPSE advocate for assigning the responsibility of generating meta-reviews to human researchers. The role of AI is to assist by identifying conflicts among reviewers’ opinions and supporting the chair in the scoring process, rather than independently assigning scores.
At present, AI for research remains in its early stages, and most systems still rely heavily on human priors and conventional research workflows. This is due to both the limitations of current models and the lack of trust from the research community. However, as model capabilities significantly improve, we may witness a paradigm shift. Systems like AlphaFold have already demonstrated that impactful scientific contributions can be made without fully replicating human research processes. In the future, AI may become an autonomous scientific agent, pursuing its own pathways of discovery, potentially letting humans learn from the model.
Discussion: The involvement of AI in manuscript writing
The application of AI in manuscript writing has been accompanied by significant controversy. As LLMs demonstrated advanced capabilities, an increasing number of researchers began adopting these systems for scholarly composition  (2404WeixinLiangMapping; 23CatherineAGaoComparing). This trend raised concerns within the academic community  (23SalvagnoCan), with scholars explicitly opposing the attribution of authorship to AI systems  (23leeCan). Despite these reservations, the substantial time efficiencies offered by this technology led researchers to gradually accept AI-assisted writing practices  (24GrudaThree; 23HuangJingshanThe; 23ChenChatGPT).This shift ultimately led to formal guidelines issued by leading academic journals  (24GanjaviPublishers; 2502ZiyangXuPatterns).
Discussion: Some areas that have not been discussed
In addition to the eight areas discussed above, there are other lines of work that also aim to support scientific research, such as reading assistance (KangHSLWH20; HeadLKFSWH21; 2303KyleLoThe), which helps researchers read academic papers; literature processing [^2] https://sdproc.org, which handles documents in various formats to provide effective data for subsequent tasks; as well as code and data generation (bauer2024comprehensive; zheng2023codegeex), which serve as a foundation for experimental validation. However, as our focus is on the core process of scientific research, we have chosen to omit these aspects from the main text.
Discussion: Unified system and domain-specific system to automate research
There is a clear distinction between unified and domain-specific AI research systems. Some efforts aim to develop general-purpose frameworks capable of supporting scientific discovery across domains (e.g., AI Scientist (2408ChrisLuThe)), while others target domain-specific challenges (e.g., AI in biology (irons2024ai)). Given the current limitations of AI capabilities, general-purpose systems have not yet fully replaced domain-specific approaches. Both directions remain valuable, but a long-term vision may favor general systems, as they hold the potential to integrate cross-disciplinary knowledge and push the boundaries of scientific understanding.
Appendix B Challenges
B.1 Hypothesis Formulation
Knowledge Synthesize
Existing paper recommendation tools predominantly rely on the metadata of existing publications to suggest related articles, which often results in a lack of user-specific targeting and insufficiently detailed presentation that hampers comprehension. Leveraging LLMs can facilitate the construction of dynamic user profiles, enabling personalized literature recommendations and enhancing the richness of the contextual information provided for each recommended article, ultimately improving the user experience. In the process of generating systematic literature reviews, our practical experience reveals that the outline generation tools often produces redundant results with insufficient hierarchical structure. Moreover, the full-text generation process is prone to hallucinations—for instance, statements may not correspond to the cited articles—a pervasive issue in large language models (2311LeiHuangA; 2402FranciscoBolaArtificial; 2404Teosusnjakautomating). This problem can be ameliorated by enhancing the foundational model capabilities or by incorporating citation tracing.
Hypothesis Generation
Most existing tools generate hypotheses by designing prompts or constructing systematic frameworks, which heavily rely on the capabilities of pre-trained models. However, these methods struggle to balance the novelty, feasibility, and validity of the hypotheses (2412RuochenLiLearning).Furthermore, our investigation reveals that many current approaches adopt novelty and feasibility as evaluation metrics; these metrics are either difficult to quantify or require manual scoring, which can introduce bias. To date, there is no unified benchmark to compare the various methods, and we believe that future research should prioritize establishing a unified metric that objectively reflects the strengths and weaknesses of different approaches.
B.2 Hypothesis Validation
Most existing scientific claim verification tools are largely confined to specific domains, exhibiting poor generalizability, which limits their practical applicability  (2305JurajVladikaScientific). Theorem proving, the scarcity of relevant data adversely affects performance improvements through training , results across different proof assistants are not directly comparable, and the lack of standardized evaluation benchmarks presents numerous challenges. Moreover, current approaches remain predominantly in the research stage and lack practical tools that facilitate interaction with researchers  (2404ZhaoyuLiA). Experiment Validation, as automatically generated experiments often suffer from a lack of methodological rigor, practical feasibility, and alignment with the original research objectives  (2410RenzeLouAAAR10). All these fields require rigorous logical reasoning, and I believe that the recent surge in advanced reasoning technologies could potentially address these issues.
B.3 Manuscript Publication
Similar to systematic literature surveys, manuscript writing is also adversely affected by hallucination issues (23AthaluriExploring; 2311LeiHuangA). Even when forced citation generation is employed, incorrect references may still be introduced (2024AljamaanReference). Furthermore, the text generated by models requires meticulous examination by researchers to avoid ethical concerns, such as plagiarism risks (23SalvagnoCan). AI-generated manuscript reviews frequently provide vague suggestions and are susceptible to biases (2405EricChamounAutomated; 24IddoDroriHuman). Additionally, during meta-review generation, models can be misled by erroneous information arising from the manuscript review process (2310SandeepKumarWhen). To address these issues, it may be necessary for the industry to establish appropriate regulations or to employ AI-based methods for detecting AI-generated papers and reviews (2111JialiangLinAutomated).
Appendix C Ability Comparison
An effective survey should not only summarize existing methods within a field but also provide comparative analyses of different approaches. However, the domain of AI for Research remains in its early stages, with many areas lacking standardized benchmarks and even established evaluation metrics. To facilitate a clearer understanding of the distinctions among various methods, we draw on existing literature (2302HyeonsuB.KangComLittee; 2402FranciscoBolaArtificial; 25LuoZimingLLM4SR; 2305JurajVladikaScientific; 2303ZonglinYangLogical; 2201XiangciLiAutomatic; 2404XiangciLiRelated; 2111JialiangLinAutomated) and adopt attribute graphs to compare representative approaches within each subfield, as illustrated in table  §3 to table  §10.
ComLittee (2302HyeonsuB.KangComLittee) - Authorship Graphs Meta data with relevant authors R, Co, Ci
ArZiGo (2410IratxePinedoArZiGo) - User Interest Meta data R
PaperWeaver (2403YoonjooLeePaperWeaver) Collected Papers Meta data with description R
22HyeonsuBKangFrom - -
Author’s social network relationships
+Reference relationship
Meta data with relevant authors R
Table 3: Research Paper Recommendation, we referred to  2302HyeonsuB.KangComLittee for comparing different methods, where R represents Paper recommender score, Co represents Co-author relationship, and Ci represents Cited author relationship.
AutoSurvey (2406YidongWangAutoSurvey) Any -
Outline Generation,
+Full-text Generation
Title & Full Content Literature Survey LLM & Human
CHIME (2407ChaoChunHsuCHIME) Biomedicine - Outline Generation Title & Full Content Hierarchical Outline Automatic Metrics
Knowledge Navigator (2408UriKatzKnowledge) Any - - Outline Generation Title & Full Content Hierarchical Outline LLM & Human
Relatedly (PalaniNDZBC23) Any - - Full-text Generation Title & Related Work Literature Survey Human
STORM (2402YijiaShaoAssisting) Any - -
Outline Generation,
+Full-text Generation
Title & Full Content Literature Survey LLM & Automatic Metrics
Table 4: Scientific Literature Review, we referred to 2402FranciscoBolaArtificial and made modifications, thereby comparing different methods.
COI (2410LongLiChain) Any - - -
Learn2Gen (2412RuochenLiLearning) Artification Intelligence - - -
MatPilot (2411ZiqiNiMatPilot) Materials Science - - -
SciAgents (2409AlirezaGhafarollahiSciAgents) Any - - - -
SciMON (2305QingyunWangSciMON) Any - - - - - -
Table 5: Hypothesis Generation, we referred to 25LuoZimingLLM4SR and made modifications, thereby comparing different methods.
MULTIVERS (2112DavidWaddenMultiVerS) Claim & scientific abstract Provided - Longformer Document Label & sentence-level rationales
SFAVEL (2309AdrinBazagaUnsupervised) Claim Pre-trained Language Model - - knowledge graph Top-K Facts & Corresponding Relevance Scores
ProToCo (2306FengzhuZengPrompt) Claim-Evidence Pair Provided - - Sentence Label
MAGIC (24WeiYuKaoMAGIC) Claim Provided - Dense Passage Retriever Sentence Label
aedFaCT (2305EnesAltuncuaedFaCT) News Article Google Search Human Document Evidence
Table 6: Scientific Claim Verification, we referred to 2305JurajVladikaScientific and made modifications, thereby comparing different methods.
IBR (QuCGDX22) - Informal -
GPT-f (2009StanislasPoluGenerative) - Formal
DT-Solver (23DT-SolverHaimingWang) Formal
POETRY (2405HaimingWangProving) - - Formal
Table 7: Theorem proving, we referred to 2303ZonglinYangLogical and made modifications, thereby comparing different methods.
AutoML-GPT (2305ShujianZhangAutoMLGPT) Artification Intelligence - - - Automated Machine Learning Task-oriented Prompts -
Chemcrow (2304AndresMBranAugmenting) Chemistry - - Chemical Task Task Description
DOLPHIN (2501YuanJiakangDolphin) Any - Automated Scientific Research Idea
MechAgents (2311BoNiMechAgents) Physics - - Mechanical Problem - -
2404ManningBenjaminSAutomated Social Science - Simulating Human - -
Method Across Stages Human Interaction Task Input Evaluation Method
AI Scientist (2408ChrisLuThe) - Full-text Generation Manuscript Template & Experimental Results & Hypothesis LLM
data-to-paper (2404TalIfarganAutonomous) Full-text Generation Experimental Results & Hypothesis -
ScholaCite (2402MartinBoyleAnnaShallow) - - Related Work Generation Title & Abstract & Citation Citation Graph Metrics
SciLit (2306NianlongGuSciLit) - Citation Generation Keywords Automatic Metrics
24GuNianlongControllable - - Citation Generation Citation Intent & Keywords Human
Table 9: Manuscript Writing, we referred to 2201XiangciLiAutomatic; 2404XiangciLiRelated and made modifications, thereby comparing different methods.
Gamma-Trans (22PanitanMuangkammuenExploiting) - - - - Peer-review Score
MARG (2401MikeDArcyMARG) - - - - Peer-review Comments
CycleResearcher (2411YixuanWengCycleResearcher) - - - Peer-review Comments & Score
PeerArg (2409PurinSukpanichnantPeerArg) - - - - - Final Decision
GLIMPSE (2406MaximeDarrinGLIMPSE) - - - - Summary of Peer-review
Table 10: Peer Review, we referred to 2111JialiangLinAutomated and made modifications, thereby comparing different methods.
[…]